### v1

@@ 0.C.2 Solver Performance on Individual Benchmarks
Alt. Loops REI.id114.ep84297 UNS 54.1 9.6 UNS 1036.7 50.9
Alt. Loops REI.id114.ep94159 SAT 0.6 6.9 SAT 22.3 42.6
Alt. Loops REI.id114.ep94581 SAT 0.8 6.9 SAT 35.9 46.2
Alt. Loops REI.id114.ep97993 SAT 0.7 7.1 SAT 139.3 48.4
Alt. Loops REI.id128.ep68334 UNS 116.6 10.1 UNS 1749.2 56.0
Alt. Loops REI.id128.ep85878 UNS 168.6 10.6 UNS 3478.2 60.4
Alt. Loops REI.id128.ep87384 SAT 0.6 6.6 SAT 308.5 57.1
Alt. Loops REI.id128.ep91042 UNS 85.5 10.3 UNS 1267.4 53.9
Alt. Loops REI.id128.ep97983 UNS 147.3 10.7 UNS 2404.6 56.7
Alt. Loops REI.id158.ep53113 SAT 25.9 10.6 SAT 16.9 34.8
Alt. Loops REI.id158.ep53499 SAT 19.4 10.6 SAT 34.1 36.8
Alt. Loops REI.id158.ep53554 SAT 12.9 10.5 SAT 17.7 39.0
Alt. Loops REI.id158.ep59378 UNS 33.3 9.3 UNS 80.8 37.6
Alt. Loops REI.id158.ep63427 UNS 28.3 9.9 UNS 68.9 38.7
Alt. Loops REI.id165.ep82822 SAT 0.7 7.0 SAT 13.2 31.6
Alt. Loops REI.id165.ep94134 SAT 0.7 7.0 SAT 14.4 33.9
Alt. Loops REI.id165.ep94522 SAT 1.0 7.1 SAT 0.2 11.2
Alt. Loops REI.id165.ep95976 SAT 0.7 7.0 SAT 30.7 34.6
Alt. Loops REI.id165.ep96974 SAT 35.8 9.7 SAT 16.8 32.5
Alt. Loops REI.id174.ep66258 SAT 16.8 10.6 SAT 0.1 5.8
Alt. Loops REI.id174.ep81057 SAT 0.6 7.0 SAT 0.3 19.6
Alt. Loops REI.id174.ep91107 SAT 25.6 10.5 SAT 49.7 43.4
Alt. Loops REI.id174.ep97933 SAT 0.7 6.8 SAT 10.6 41.8
Alt. Loops REI.id174.ep98455 SAT 21.5 10.9 SAT 0.2 13.0
Alt. Loops REI.id176.ep92206 UNS 635.3 11.0 UNK 3601.0 66.0
Alt. Loops REI.id176.ep95516 UNS 375.7 10.6 UNK 3601.0 61.5
Alt. Loops REI.id176.ep96138 UNS 340.2 10.6 UNK 3601.0 63.1
Alt. Loops REI.id176.ep98430 UNS 548.5 10.9 UNK 3601.0 64.6
Alt. Loops REI.id176.ep99427 UNS 536.1 10.7 UNK 3601.0 64.2
Alt. Loops REI.id180.ep58329 SAT 0.5 6.7 SAT 0.3 15.8
Alt. Loops REI.id180.ep58461 SAT 0.7 6.7 SAT 0.2 14.3
Alt. Loops REI.id180.ep61293 SAT 0.8 7.0 SAT 0.6 28.1
Alt. Loops REI.id180.ep69795 SAT 0.7 6.9 SAT 0.2 9.6
Alt. Loops REI.id180.ep77624 SAT 0.7 7.2 SAT 0.1 9.3
Alt. Loops REI.id196.ep41633 SAT 0.9 7.0 SAT 0.2 12.0
Alt. Loops REI.id196.ep43251 SAT 0.8 7.0 SAT 0.1 9.3
Alt. Loops REI.id196.ep73462 SAT 0.7 6.9 SAT 0.2 12.8
Alt. Loops REI.id196.ep86142 SAT 0.8 7.1 SAT 0.1 8.3
Alt. Loops REI.id196.ep99092 SAT 0.6 6.9 SAT 24.0 34.0
Alt. Loops REI.id201.ep39282 SAT 17.4 10.1 SAT 4.0 39.8
Alt. Loops REI.id201.ep48186 SAT 0.7 7.0 SAT 51.4 43.1
Alt. Loops REI.id201.ep48219 SAT 0.6 6.6 SAT 0.2 10.6
Alt. Loops REI.id201.ep56910 SAT 0.8 7.1 SAT 28.7 34.9
Alt. Loops REI.id201.ep88825 SAT 0.6 6.5 SAT 49.6 45.4
Alt. Loops REI.id215.ep96277 SAT 0.3 6.6 SAT 1.7 41.1
Alt. Loops REI.id215.ep97333 SAT 0.4 6.3 SAT 4.4 40.8
Alt. Loops REI.id215.ep97355 SAT 0.3 6.6 SAT 0.6 32.9
Alt. Loops REI.id215.ep97421 SAT 0.7 6.6 SAT 1763.0 60.1
Alt. Loops REI.id215.ep98767 SAT 0.6 6.9 SAT 2.1 35.0
Alt. Loops REI.id234.ep70556 SAT 0.7 6.8 SAT 0.1 10.8
Alt. Loops REI.id234.ep71035 SAT 0.6 6.6 SAT 0.3 24.4
Alt. Loops REI.id234.ep71965 SAT 0.7 7.0 SAT 0.2 11.8
Alt. Loops REI.id234.ep83703 SAT 0.5 6.9 SAT 0.3 22.4
Alt. Loops REI.id234.ep83780 SAT 0.7 8.0 SAT 1.0 33.8
Alt. Loops REI.id239.ep71386 UNS 197.2 10.3 UNS 1691.7 55.5
Alt. Loops REI.id239.ep73100 UNS 186.9 10.6 UNS 1110.3 60.0
Alt. Loops REI.id239.ep73504 SAT 59.0 11.1 SAT 3.4 51.2
Alt. Loops REI.id239.ep79112 UNS 190.0 10.4 UNS 1643.7 56.7
Alt. Loops REI.id239.ep90118 SAT 30.8 9.9 SAT 0.4 18.0
Alt. Loops REI.id240.ep85050 UNS 40.3 10.0 UNS 308.5 47.0
Alt. Loops REI.id240.ep89317 UNS 32.5 10.7 UNS 152.7 42.7
Alt. Loops REI.id240.ep91714 UNS 26.8 9.8 UNS 177.1 42.5
Alt. Loops REI.id240.ep98115 UNS 40.2 10.7 UNS 282.4 47.6
Alt. Loops REI.id240.ep98891 UNS 31.9 10.2 UNS 76.3 45.4
Alt. Loops REI.id267.ep91426 UNS 0.9 6.5 UNS 0.7 14.5
Alt. Loops REI.id267.ep95713 UNS 1.1 6.5 UNS 2.5 16.7
Alt. Loops REI.id267.ep96300 UNS 0.9 6.8 UNS 1.0 14.8
Alt. Loops REI.id267.ep96333 UNS 1.4 7.0 UNS 0.5 12.4
Alt. Loops REI.id267.ep96426 UNS 1.5 6.9 UNS 1.5 15.5
Alt. Loops REI.id286.ep89988 UNS 38.1 9.1 UNS 41.8 35.9
Alt. Loops REI.id286.ep89999 UNS 38.4 9.1 UNS 41.7 35.8
Alt. Loops REI.id286.ep90032 UNS 32.5 9.7 UNS 59.7 37.9
Alt. Loops REI.id286.ep94925 UNS 34.0 9.0 UNS 56.8 38.1
Alt. Loops REI.id286.ep95719 UNS 27.6 9.4 UNS 63.4 38.2
Alt. Loops REI.id296.ep74854 UNS 127.3 11.1 UNS 697.8 51.9
Alt. Loops REI.id296.ep74969 UNS 108.8 10.0 UNS 559.8 55.9
Alt. Loops REI.id296.ep76414 UNS 113.8 10.3 UNS 500.1 51.7
Alt. Loops REI.id296.ep77867 UNS 116.8 10.0 UNS 131.1 48.4
Alt. Loops REI.id296.ep79125 UNS 111.0 10.4 UNS 274.0 46.4
Alt. Loops REI.id298.ep75768 UNS 167.3 9.8 UNK 3601.0 66.3
Alt. Loops REI.id298.ep77187 SAT 0.8 6.9 SAT 0.4 36.5
Alt. Loops REI.id298.ep79143 UNS 162.0 10.3 UNK 3601.0 70.0
Alt. Loops REI.id298.ep85589 UNS 198.8 11.0 UNK 3601.0 65.9
Alt. Loops REI.id298.ep97154 SAT 0.6 7.0 SAT 0.1 8.0
Alt. Loops REI.id299.ep69103 UNS 122.0 11.0 UNS 479.8 49.8
Alt. Loops REI.id299.ep69400 UNS 98.1 10.9 UNS 361.7 47.0
Alt. Loops REI.id299.ep81175 UNS 79.2 10.6 UNS 284.3 45.7
Alt. Loops REI.id299.ep92705 UNS 90.0 10.6 UNS 739.8 50.7
Alt. Loops REI.id299.ep99426 UNS 49.8 10.3 UNS 569.1 49.5
Alt. Loops REI.id303.ep68520 UNS 402.2 12.0 UNK 3601.0 64.6
Alt. Loops REI.id303.ep69708 SAT 1.7 7.6 SAT 987.6 61.3
Alt. Loops REI.id303.ep94994 SAT 0.7 6.6 SAT 72.8 58.6
Alt. Loops REI.id303.ep95038 SAT 0.9 7.1 SAT 4.4 52.3
Alt. Loops REI.id303.ep95137 SAT 1.4 7.2 SAT 1399.1 72.3
Alt. Loops REI.id308.ep94787 UNS 39.2 9.8 UNS 95.0 46.1
Alt. Loops REI.id308.ep97397 SAT 0.7 7.3 SAT 0.2 14.4
Alt. Loops REI.id308.ep98363 SAT 0.6 7.3 SAT 125.7 45.1
Alt. Loops REI.id308.ep98830 UNS 47.3 9.8 UNS 207.0 46.9
Alt. Loops REI.id308.ep99303 SAT 1.0 6.7 SAT 6.1 35.9
Alt. Loops REI.id318.ep75036 SAT 0.4 6.6 SAT 14.4 64.8
Alt. Loops REI.id318.ep85341 SAT 50.0 10.4 SAT 164.6 57.5
Alt. Loops REI.id318.ep89536 SAT 0.8 7.0 SAT 27.3 55.8
Alt. Loops REI.id318.ep94480 SAT 0.7 6.9 SAT 0.9 37.7
Alt. Loops REI.id318.ep96308 SAT 0.7 7.0 SAT 0.4 21.7
Alt. Loops REI.id319.ep62661 SAT 0.8 6.9 SAT 71.0 54.8
Alt. Loops REI.id319.ep84434 SAT 0.6 7.0 SAT 2437.7 67.5
Alt. Loops REI.id319.ep84478 SAT 0.8 7.0 SAT 0.3 14.3
Alt. Loops REI.id319.ep86630 SAT 0.8 7.0 SAT 573.1 62.2
Alt. Loops REI.id319.ep89649 SAT 0.7 6.7 SAT 0.3 12.9
Alt. Loops REI.id321.ep55647 UNS 9.1 9.2 UNS 19.9 33.0
Alt. Loops REI.id321.ep81083 UNS 34.1 8.8 UNS 404.1 53.6
Alt. Loops REI.id321.ep86834 UNS 25.6 9.2 UNS 712.3 62.0
Alt. Loops REI.id321.ep92994 UNS 48.9 9.9 UNS 389.5 54.3
Alt. Loops REI.id321.ep93439 UNS 24.8 9.9 UNS 686.8 58.2
Alt. Loops REI.id343.ep55239 UNS 83.2 9.2 UNS 149.7 47.1
Alt. Loops REI.id343.ep63040 UNS 77.5 9.8 UNS 168.2 45.4
Alt. Loops REI.id343.ep63051 UNS 76.7 10.3 UNS 166.6 45.5
Alt. Loops REI.id343.ep73171 UNS 160.2 10.2 UNS 1022.7 54.2
Alt. Loops REI.id343.ep96195 UNS 241.5 9.9 UNS 3113.7 57.7
Alt. Loops REI.id35.ep61486 SAT 0.4 6.3 SAT 22.1 39.1
Alt. Loops REI.id35.ep61788 SAT 0.4 6.5 SAT 0.2 10.9
Alt. Loops REI.id35.ep81067 SAT 0.6 7.2 SAT 0.2 12.3
Alt. Loops REI.id35.ep85842 SAT 0.6 6.6 SAT 0.2 10.4
Alt. Loops REI.id35.ep94624 SAT 0.3 6.4 SAT 0.1 7.1
Alt. Loops REI.id352.ep49376 UNS 8.6 9.1 UNS 58.8 39.3
Alt. Loops REI.id352.ep51510 UNS 12.2 8.6 UNS 169.9 43.4
Alt. Loops REI.id352.ep68171 UNS 13.4 9.6 UNS 87.1 39.8
Alt. Loops REI.id352.ep69872 UNS 10.8 9.4 UNS 68.1 38.3
Alt. Loops REI.id352.ep70312 UNS 12.6 9.2 UNS 24.0 37.8
Alt. Loops REI.id379.ep71969 SAT 3.8 9.8 SAT 13.7 43.2
Alt. Loops REI.id379.ep80359 SAT 0.6 6.7 SAT 0.2 11.7
Alt. Loops REI.id379.ep87708 UNS 214.8 11.1 UNS 2425.8 57.3
Alt. Loops REI.id379.ep96909 SAT 1.9 8.7 SAT 180.9 69.4
Alt. Loops REI.id379.ep97133 SAT 0.5 6.7 SAT 0.7 36.4
Alt. Loops REI.id381.ep85470 SAT 58.1 12.1 UNK 3601.0 64.0
Alt. Loops REI.id381.ep85558 SAT 0.7 6.7 SAT 1140.9 62.2
Alt. Loops REI.id381.ep92541 UNS 321.6 10.9 UNK 3600.9 59.8
Alt. Loops REI.id381.ep92994 SAT 1.6 7.4 UNK 3601.0 64.2
Alt. Loops REI.id381.ep93562 UNS 315.7 11.5 UNK 3601.0 64.3
Alt. Loops REI.id393.ep88789 SAT 18.4 11.9 SAT 0.3 17.0
Alt. Loops REI.id393.ep90550 SAT 0.8 7.2 SAT 64.9 56.5
Alt. Loops REI.id393.ep97111 SAT 19.3 11.6 SAT 0.1 10.0
Alt. Loops REI.id393.ep97199 SAT 23.3 11.3 SAT 0.2 15.2
Alt. Loops REI.id393.ep97221 SAT 23.8 10.4 SAT 0.2 14.3
Alt. Loops REI.id399.ep62611 UNS 45.5 9.7 UNS 490.9 56.3
Alt. Loops REI.id399.ep69315 UNS 56.0 10.4 UNS 3342.7 68.6
Alt. Loops REI.id399.ep85537 UNS 76.4 9.5 UNS 2220.3 60.8
Alt. Loops REI.id399.ep92145 UNS 69.7 10.6 UNS 1475.9 64.6
Alt. Loops REI.id399.ep96449 SAT 1.1 7.1 SAT 140.4 64.8
Alt. Loops REI.id418.ep91524 SAT 0.8 6.9 SAT 5.3 53.1
Alt. Loops REI.id418.ep92561 SAT 1.2 7.1 SAT 19.8 51.7
Alt. Loops REI.id418.ep92671 SAT 1.5 7.4 SAT 30.0 49.7
Alt. Loops REI.id418.ep99126 SAT 96.6 11.0 SAT 134.2 51.4
Alt. Loops REI.id418.ep99947 SAT 0.7 7.1 SAT 13.4 51.0
Alt. Loops REI.id425.ep69300 SAT 0.7 6.6 SAT 0.8 37.0
Alt. Loops REI.id425.ep83340 SAT 0.3 6.5 SAT 0.3 14.0
Alt. Loops REI.id425.ep93745 UNS 166.7 11.0 UNS 1299.1 51.4
Alt. Loops REI.id425.ep96981 SAT 0.7 6.9 SAT 0.1 8.0
Alt. Loops REI.id425.ep98177 SAT 0.8 7.1 SAT 0.1 11.8
Alt. Loops REI.id444.ep77347 UNS 90.0 10.3 UNS 501.8 50.7
Alt. Loops REI.id444.ep78487 UNS 89.7 11.1 UNS 1094.5 51.6
Alt. Loops REI.id444.ep87148 UNS 71.6 9.8 UNS 842.5 51.1
Alt. Loops REI.id444.ep92635 UNS 96.3 10.8 UNK 3601.0 48.8
Alt. Loops REI.id444.ep96160 UNS 95.2 10.1 UNS 1262.5 51.4
Alt. Loops REI.id457.ep46821 SAT 9.4 8.8 SAT 0.1 8.8
Alt. Loops REI.id457.ep69348 SAT 0.4 6.5 SAT 0.2 10.6
Alt. Loops REI.id457.ep78403 SAT 0.7 7.1 SAT 0.1 7.3
Alt. Loops REI.id457.ep79411 SAT 0.6 6.7 SAT 0.2 12.7
Alt. Loops REI.id457.ep81404 SAT 0.5 6.8 SAT 0.1 8.5
Alt. Loops REI.id47.ep79665 SAT 1.1 7.0 SAT 0.3 19.2
Alt. Loops REI.id47.ep92041 SAT 0.6 6.7 SAT 0.1 10.0
Alt. Loops REI.id47.ep96122 SAT 0.3 6.5 SAT 0.4 21.0
Alt. Loops REI.id47.ep99600 SAT 0.2 6.5 SAT 0.3 18.8
[…]

### v2

@@ 0.C.2 Solver Performance on Individual Benchmarks
Alt. Loops REI.id114.ep84297 UNSAT 55.7 10.6 UNSAT 1036.7 50.9
Alt. Loops REI.id114.ep94159 SAT 0.6 6.7 SAT 22.3 42.6
Alt. Loops REI.id114.ep94581 SAT 0.6 6.7 SAT 35.9 46.2
Alt. Loops REI.id114.ep97993 SAT 0.5 6.7 SAT 139.3 48.4
Alt. Loops REI.id128.ep68334 UNSAT 111.0 9.6 UNSAT 1749.2 56.0
Alt. Loops REI.id128.ep85878 UNSAT 171.1 10.2 UNSAT 3478.2 60.4
Alt. Loops REI.id128.ep87384 SAT 0.7 6.3 SAT 308.5 57.1
Alt. Loops REI.id128.ep91042 UNSAT 94.9 9.2 UNSAT 1267.4 53.9
Alt. Loops REI.id128.ep97983 UNSAT 127.2 10.0 UNSAT 2404.6 56.7
Alt. Loops REI.id158.ep53113 SAT 70.9 9.7 SAT 16.9 34.8
Alt. Loops REI.id158.ep53499 SAT 14.6 8.6 SAT 34.1 36.8
Alt. Loops REI.id158.ep53554 SAT 10.4 8.9 SAT 17.7 39.0
Alt. Loops REI.id158.ep59378 UNSAT 32.0 9.8 UNSAT 80.8 37.6
Alt. Loops REI.id158.ep63427 UNSAT 23.7 9.4 UNSAT 68.9 38.7
Alt. Loops REI.id165.ep82822 SAT 0.7 6.8 SAT 13.2 31.6
Alt. Loops REI.id165.ep94134 SAT 0.7 6.7 SAT 14.4 33.9
Alt. Loops REI.id165.ep94522 SAT 0.5 6.5 SAT 0.2 11.2
Alt. Loops REI.id165.ep95976 SAT 0.5 6.8 SAT 30.7 34.6
Alt. Loops REI.id165.ep96974 SAT 24.0 8.3 SAT 16.8 32.5
Alt. Loops REI.id174.ep66258 SAT 10.5 9.6 SAT 0.1 5.8
Alt. Loops REI.id174.ep81057 SAT 0.6 6.5 SAT 0.3 19.6
Alt. Loops REI.id174.ep91107 SAT 18.0 9.5 SAT 49.7 43.4
Alt. Loops REI.id174.ep97933 SAT 0.7 6.7 SAT 10.6 41.8
Alt. Loops REI.id174.ep98455 SAT 25.5 10.3 SAT 0.2 13.0
Alt. Loops REI.id176.ep92206 UNSAT 507.3 10.8 UNK 3601.0 66.0
Alt. Loops REI.id176.ep95516 UNSAT 438.3 10.4 UNK 3601.0 61.5
Alt. Loops REI.id176.ep96138 UNSAT 364.7 10.1 UNK 3601.0 63.1
Alt. Loops REI.id176.ep98430 UNSAT 816.4 10.8 UNK 3601.0 64.6
Alt. Loops REI.id176.ep99427 UNSAT 410.6 10.3 UNK 3601.0 64.2
Alt. Loops REI.id180.ep58329 SAT 0.5 6.2 SAT 0.3 15.8
Alt. Loops REI.id180.ep58461 SAT 0.5 6.4 SAT 0.2 14.3
Alt. Loops REI.id180.ep61293 SAT 0.6 6.4 SAT 0.6 28.1
Alt. Loops REI.id180.ep69795 SAT 0.5 6.4 SAT 0.2 9.6
Alt. Loops REI.id180.ep77624 SAT 0.4 6.6 SAT 0.1 9.3
Alt. Loops REI.id196.ep41633 SAT 0.8 6.6 SAT 0.2 12.0
Alt. Loops REI.id196.ep43251 SAT 0.3 6.4 SAT 0.1 9.3
Alt. Loops REI.id196.ep73462 SAT 0.7 6.7 SAT 0.2 12.8
Alt. Loops REI.id196.ep86142 SAT 0.3 6.4 SAT 0.1 8.3
Alt. Loops REI.id196.ep99092 SAT 0.4 6.2 SAT 24.0 34.0
Alt. Loops REI.id201.ep39282 SAT 21.5 9.6 SAT 4.0 39.8
Alt. Loops REI.id201.ep48186 SAT 0.6 6.2 SAT 51.4 43.1
Alt. Loops REI.id201.ep48219 SAT 0.6 6.4 SAT 0.2 10.6
Alt. Loops REI.id201.ep56910 SAT 0.4 6.4 SAT 28.7 34.9
Alt. Loops REI.id201.ep88825 SAT 0.2 6.0 SAT 49.6 45.4
Alt. Loops REI.id215.ep96277 SAT 0.5 6.6 SAT 1.7 41.1
Alt. Loops REI.id215.ep97333 SAT 0.3 6.4 SAT 4.4 40.8
Alt. Loops REI.id215.ep97355 SAT 0.6 6.4 SAT 0.6 32.9
Alt. Loops REI.id215.ep97421 SAT 0.5 6.9 SAT 1763.0 60.1
Alt. Loops REI.id215.ep98767 SAT 0.5 6.5 SAT 2.1 35.0
Alt. Loops REI.id234.ep70556 SAT 0.6 6.5 SAT 0.1 10.8
Alt. Loops REI.id234.ep71035 SAT 0.3 6.2 SAT 0.3 24.4
Alt. Loops REI.id234.ep71965 SAT 0.6 7.1 SAT 0.2 11.8
Alt. Loops REI.id234.ep83703 SAT 0.7 6.7 SAT 0.3 22.4
Alt. Loops REI.id234.ep83780 SAT 0.4 6.4 SAT 1.0 33.8
Alt. Loops REI.id239.ep71386 UNSAT 209.9 10.2 UNSAT 1691.7 55.5
Alt. Loops REI.id239.ep73100 UNSAT 210.3 10.4 UNSAT 1110.3 60.0
Alt. Loops REI.id239.ep73504 SAT 6.7 9.1 SAT 3.4 51.2
Alt. Loops REI.id239.ep79112 UNSAT 193.2 9.9 UNSAT 1643.7 56.7
Alt. Loops REI.id239.ep90118 SAT 35.8 9.3 SAT 0.4 18.0
Alt. Loops REI.id240.ep85050 UNSAT 37.2 9.7 UNSAT 308.5 47.0
Alt. Loops REI.id240.ep89317 UNSAT 26.5 9.4 UNSAT 152.7 42.7
Alt. Loops REI.id240.ep91714 UNSAT 28.2 9.9 UNSAT 177.1 42.5
Alt. Loops REI.id240.ep98115 UNSAT 40.2 10.2 UNSAT 282.4 47.6
Alt. Loops REI.id240.ep98891 UNSAT 28.1 9.6 UNSAT 76.3 45.4
Alt. Loops REI.id267.ep91426 UNSAT 0.7 6.2 UNSAT 0.7 14.5
Alt. Loops REI.id267.ep95713 UNSAT 1.7 6.4 UNSAT 2.5 16.7
Alt. Loops REI.id267.ep96300 UNSAT 1.7 6.9 UNSAT 1.0 14.8
Alt. Loops REI.id267.ep96333 UNSAT 1.1 6.2 UNSAT 0.5 12.4
Alt. Loops REI.id267.ep96426 UNSAT 1.3 6.0 UNSAT 1.5 15.5
Alt. Loops REI.id286.ep89988 UNSAT 32.9 8.5 UNSAT 41.8 35.9
Alt. Loops REI.id286.ep89999 UNSAT 32.4 8.5 UNSAT 41.7 35.8
Alt. Loops REI.id286.ep90032 UNSAT 39.1 9.0 UNSAT 59.7 37.9
Alt. Loops REI.id286.ep94925 UNSAT 35.1 9.2 UNSAT 56.8 38.1
Alt. Loops REI.id286.ep95719 UNSAT 50.8 9.0 UNSAT 63.4 38.2
Alt. Loops REI.id296.ep74854 UNSAT 140.7 9.7 UNSAT 697.8 51.9
Alt. Loops REI.id296.ep74969 UNSAT 119.7 9.6 UNSAT 559.8 55.9
Alt. Loops REI.id296.ep76414 UNSAT 127.1 10.1 UNSAT 500.1 51.7
Alt. Loops REI.id296.ep77867 UNSAT 100.3 9.4 UNSAT 131.1 48.4
Alt. Loops REI.id296.ep79125 UNSAT 133.1 9.9 UNSAT 274.0 46.4
Alt. Loops REI.id298.ep75768 UNSAT 240.6 10.9 UNK 3601.0 66.3
Alt. Loops REI.id298.ep77187 SAT 0.8 6.8 SAT 0.4 36.5
Alt. Loops REI.id298.ep79143 UNSAT 167.6 10.9 UNK 3601.0 70.0
Alt. Loops REI.id298.ep85589 UNSAT 175.5 10.8 UNK 3601.0 65.9
Alt. Loops REI.id298.ep97154 SAT 0.5 6.8 SAT 0.1 8.0
Alt. Loops REI.id299.ep69103 UNSAT 57.5 11.2 UNSAT 479.8 49.8
Alt. Loops REI.id299.ep69400 UNSAT 49.6 10.7 UNSAT 361.7 47.0
Alt. Loops REI.id299.ep81175 UNSAT 40.3 9.7 UNSAT 284.3 45.7
Alt. Loops REI.id299.ep92705 UNSAT 111.3 10.7 UNSAT 739.8 50.7
Alt. Loops REI.id299.ep99426 UNSAT 69.2 9.8 UNSAT 569.1 49.5
Alt. Loops REI.id303.ep68520 UNSAT 243.1 11.6 UNK 3601.0 64.6
Alt. Loops REI.id303.ep69708 SAT 1.6 8.6 SAT 987.6 61.3
Alt. Loops REI.id303.ep94994 SAT 0.6 6.6 SAT 72.8 58.6
Alt. Loops REI.id303.ep95038 SAT 0.6 6.4 SAT 4.4 52.3
Alt. Loops REI.id303.ep95137 SAT 0.4 6.4 SAT 1399.1 72.3
Alt. Loops REI.id308.ep94787 UNSAT 19.9 8.7 UNSAT 95.0 46.1
Alt. Loops REI.id308.ep97397 SAT 0.5 6.7 SAT 0.2 14.4
Alt. Loops REI.id308.ep98363 SAT 0.7 6.6 SAT 125.7 45.1
Alt. Loops REI.id308.ep98830 UNSAT 52.5 10.1 UNSAT 207.0 46.9
Alt. Loops REI.id308.ep99303 SAT 0.8 6.9 SAT 6.1 35.9
Alt. Loops REI.id318.ep75036 SAT 0.6 6.3 SAT 14.4 64.8
Alt. Loops REI.id318.ep85341 SAT 96.4 10.7 SAT 164.6 57.5
Alt. Loops REI.id318.ep89536 SAT 0.7 6.5 SAT 27.3 55.8
Alt. Loops REI.id318.ep94480 SAT 0.5 6.5 SAT 0.9 37.7
Alt. Loops REI.id318.ep96308 SAT 0.6 6.7 SAT 0.4 21.7
Alt. Loops REI.id319.ep62661 SAT 0.4 6.5 SAT 71.0 54.8
Alt. Loops REI.id319.ep84434 SAT 0.6 6.9 SAT 2437.7 67.5
Alt. Loops REI.id319.ep84478 SAT 0.3 6.4 SAT 0.3 14.3
Alt. Loops REI.id319.ep86630 SAT 0.5 6.2 SAT 573.1 62.2
Alt. Loops REI.id319.ep89649 SAT 0.3 6.6 SAT 0.3 12.9
Alt. Loops REI.id321.ep55647 UNSAT 9.7 8.5 UNSAT 19.9 33.0
Alt. Loops REI.id321.ep81083 UNSAT 18.2 8.5 UNSAT 404.1 53.6
Alt. Loops REI.id321.ep86834 UNSAT 28.6 9.3 UNSAT 712.3 62.0
Alt. Loops REI.id321.ep92994 UNSAT 25.0 9.5 UNSAT 389.5 54.3
Alt. Loops REI.id321.ep93439 UNSAT 40.2 8.9 UNSAT 686.8 58.2
Alt. Loops REI.id343.ep55239 UNSAT 62.6 9.2 UNSAT 149.7 47.1
Alt. Loops REI.id343.ep63040 UNSAT 83.4 9.1 UNSAT 168.2 45.4
Alt. Loops REI.id343.ep63051 UNSAT 83.2 9.1 UNSAT 166.6 45.5
Alt. Loops REI.id343.ep73171 UNSAT 156.1 9.7 UNSAT 1022.7 54.2
Alt. Loops REI.id343.ep96195 UNSAT 235.0 9.3 UNSAT 3113.7 57.7
Alt. Loops REI.id35.ep61486 SAT 0.4 6.4 SAT 22.1 39.1
Alt. Loops REI.id35.ep61788 SAT 0.5 6.7 SAT 0.2 10.9
Alt. Loops REI.id35.ep81067 SAT 0.3 6.3 SAT 0.2 12.3
Alt. Loops REI.id35.ep85842 SAT 0.2 6.1 SAT 0.2 10.4
Alt. Loops REI.id35.ep94624 SAT 0.4 6.4 SAT 0.1 7.1
Alt. Loops REI.id352.ep49376 UNSAT 7.6 8.2 UNSAT 58.8 39.3
Alt. Loops REI.id352.ep51510 UNSAT 8.0 9.0 UNSAT 169.9 43.4
Alt. Loops REI.id352.ep68171 UNSAT 12.1 9.2 UNSAT 87.1 39.8
Alt. Loops REI.id352.ep69872 UNSAT 9.5 8.8 UNSAT 68.1 38.3
Alt. Loops REI.id352.ep70312 UNSAT 8.1 8.2 UNSAT 24.0 37.8
Alt. Loops REI.id379.ep71969 SAT 1.4 7.9 SAT 13.7 43.2
Alt. Loops REI.id379.ep80359 SAT 0.7 6.8 SAT 0.2 11.7
Alt. Loops REI.id379.ep87708 UNSAT 212.5 10.6 UNSAT 2425.8 57.3
Alt. Loops REI.id379.ep96909 SAT 32.1 10.7 SAT 180.9 69.4
Alt. Loops REI.id379.ep97133 SAT 1.2 7.2 SAT 0.7 36.4
Alt. Loops REI.id381.ep85470 SAT 1.9 9.2 UNK 3601.0 64.0
Alt. Loops REI.id381.ep85558 SAT 1.7 8.8 SAT 1140.9 62.2
Alt. Loops REI.id381.ep92541 UNSAT 250.5 11.2 UNK 3600.9 59.8
Alt. Loops REI.id381.ep92994 SAT 1.3 7.3 UNK 3601.0 64.2
Alt. Loops REI.id381.ep93562 UNSAT 281.6 10.7 UNK 3601.0 64.3
Alt. Loops REI.id393.ep88789 SAT 25.5 11.0 SAT 0.3 17.0
Alt. Loops REI.id393.ep90550 SAT 0.5 6.4 SAT 64.9 56.5
Alt. Loops REI.id393.ep97111 SAT 31.6 10.0 SAT 0.1 10.0
Alt. Loops REI.id393.ep97199 SAT 43.0 10.6 SAT 0.2 15.2
Alt. Loops REI.id393.ep97221 SAT 37.3 11.0 SAT 0.2 14.3
Alt. Loops REI.id399.ep62611 UNSAT 42.9 9.2 UNSAT 490.9 56.3
Alt. Loops REI.id399.ep69315 UNSAT 43.9 9.8 UNSAT 3342.7 68.6
Alt. Loops REI.id399.ep85537 UNSAT 91.9 9.5 UNSAT 2220.3 60.8
Alt. Loops REI.id399.ep92145 UNSAT 47.1 8.8 UNSAT 1475.9 64.6
Alt. Loops REI.id399.ep96449 SAT 1.5 7.1 SAT 140.4 64.8
Alt. Loops REI.id418.ep91524 SAT 0.4 6.2 SAT 5.3 53.1
Alt. Loops REI.id418.ep92561 SAT 0.6 6.6 SAT 19.8 51.7
Alt. Loops REI.id418.ep92671 SAT 0.5 6.4 SAT 30.0 49.7
Alt. Loops REI.id418.ep99126 SAT 0.9 6.4 SAT 134.2 51.4
Alt. Loops REI.id418.ep99947 SAT 0.6 6.5 SAT 13.4 51.0
Alt. Loops REI.id425.ep69300 SAT 0.6 6.7 SAT 0.8 37.0
Alt. Loops REI.id425.ep83340 SAT 0.6 6.9 SAT 0.3 14.0
Alt. Loops REI.id425.ep93745 UNSAT 113.7 10.0 UNSAT 1299.1 51.4
Alt. Loops REI.id425.ep96981 SAT 0.5 6.6 SAT 0.1 8.0
Alt. Loops REI.id425.ep98177 SAT 0.5 6.6 SAT 0.1 11.8
Alt. Loops REI.id444.ep77347 UNSAT 90.7 10.7 UNSAT 501.8 50.7
Alt. Loops REI.id444.ep78487 UNSAT 120.5 10.2 UNSAT 1094.5 51.6
Alt. Loops REI.id444.ep87148 UNSAT 78.3 10.0 UNSAT 842.5 51.1
Alt. Loops REI.id444.ep92635 UNSAT 105.6 10.1 UNK 3601.0 48.8
Alt. Loops REI.id444.ep96160 UNSAT 89.3 9.7 UNSAT 1262.5 51.4
Alt. Loops REI.id457.ep46821 SAT 11.0 9.3 SAT 0.1 8.8
Alt. Loops REI.id457.ep69348 SAT 0.5 6.9 SAT 0.2 10.6
Alt. Loops REI.id457.ep78403 SAT 0.3 6.4 SAT 0.1 7.3
Alt. Loops REI.id457.ep79411 SAT 0.5 6.3 SAT 0.2 12.7
Alt. Loops REI.id457.ep81404 SAT 0.6 6.5 SAT 0.1 8.5
Alt. Loops REI.id47.ep79665 SAT 0.8 6.6 SAT 0.3 19.2
Alt. Loops REI.id47.ep92041 SAT 0.3 6.3 SAT 0.1 10.0
Alt. Loops REI.id47.ep96122 SAT 0.7 6.6 SAT 0.4 21.0
Alt. Loops REI.id47.ep99600 SAT 0.6 6.4 SAT 0.3 18.8
[…]